A reaction has a rate constant of 1.26×10−4 s−1 at 26  ∘C and 0.229 s−1 at 77  ∘C .

A) Determine the activation barrier for the reaction. ( kJ/mol )

B) What is the value of the rate constant at 16  ∘C ? (s^-1)
Answer #1

Given that

K1 = 1.26*10-4 s-1

temperature (T1) = 26oC = 299.15K

k2 = 0.229 s-1

Temperature (T2) = 77oC = 77+ 273.15 = 350.15 K

Gas constant (R) = 8.314 J/mol.K

As we know

ln (k2/k1) = (Ea /R)(1/T1 - 1/T2)

ln ( 0.229 s-1 / 1.26*10-4 s-1) = (Ea / 8.314 J/mol.K) (1/299.15K - 1/350.15K)

Ea = Activation energy = 128157.89 J = 128.2 kJ /mol

(B)Again if T2 = 16oC = 16+273.15 = 289.15 K and k2 = ?

ln (k2/k1) = (Ea /R)(1/T1 - 1/T2)

ln ( k2 / 1.26*10-4 s-1) = (128157.89 J / 8.314 J/mol.K) (1/299.15K - 1/289.15K)

k2 = 2.12 *10-5 s-1